OpenStudy (anonymous):

y-3=m(x-6) how do we get it into y=mx+b form plaese helppppppp

jimthompson5910 (jim_thompson5910):

y-3=m(x-6) y-3=mx-6m y=mx-6m+3 The equation is now in y = mx+b form where m is itself and b = -6m+3